OpenStudy (anonymous):

which of the following fractions is equivalent to the fraction 6/16? A.) 4/14 B.) 1/10 C.) 3/8 D.) 2/8

OpenStudy (oaktree):

So we want to simplify the fraction. It helps to write out the numerator and the denominator as products - like prime factorize them both. What are the prime factorizations of 6 and 16?

OpenStudy (zale101):

Do u know how to prime factorize?

OpenStudy (zale101):

for example: The prime factors of 9 are 3 and 3

OpenStudy (zale101):

What are the prime factorization of 6 and 16?

OpenStudy (anonymous):

c)

OpenStudy (anonymous):

3/8

OpenStudy (phi):

to simplify 6/16 you should write 6 as 2*3 notice 16 can be divide by 2: 2*8 but 8 can be divided by 2: 2*2*4, and so on you get 16=2*2*2*2 the fraction \[ \frac{6}{16}= \frac{2 \cdot 3}{ 2\cdot 2\cdot 2\cdot 2} \] now look for the same number in the top and bottom. They "cancel out", which means they = 1, and 1 times anything does not change the answer \[ \frac{6}{16}= \frac{\cancel{2} \cdot 3}{ \cancel{2}\cdot 2\cdot 2\cdot 2} \] now multiply the bottom 2's to get the final answer. Of course, in this case, we know 16 (or 8) is not divisible by 3 so we could have been smart an written the bottom as 2*8 btw, to do these problems you should learn the "divisibility rules" http://math.about.com/library/bldivide.htm and memorize the first few prime numbers http://primes.utm.edu/lists/small/1000.txt (1000 is overkill, but the first line would be great)

OpenStudy (anonymous):

C 3/8